OpenStudy (anonymous):

A line passes through the points (-10,-4) and (-1,2) . What is the y-intercept of the line? HintX

OpenStudy (anonymous):

OK, let's make an equation: Slope = 2 - (-4) / -1 - (-10) = 6 / 9 = 2/3. So, the equation is 2/3x + k = y. Plug in x and y and it's 2/3 * -1 + k = 2. -2/3 + k = 2. k = 2 2/3 or 8/3.

OpenStudy (anonymous):

So will it be {0,8/3) or (8/3,0)

OpenStudy (anonymous):

The point is (0, 8/3); all y intercepts are in the form (0, k)